A bag contains four marbles: one red, one black, one white, and one orange. Two marbles are to be selected at random from the bag, with replacement. List the sample space. Find the probability that the black marble is chosen twice.
Answer by ewatrrr(10682)

 
Hi,
A bag contains four marbles: one red, one black, one white, and one orange.
Two marbles are to be selected at random from the bag, with replacement.
Sample Space: RR,BB,WW,OO, RB,BR,RW,WR,RO,OR,BW,WB,B0,OB, WO,OW
P(2B) = 1/4*1/4 = 1/16